How to solve this simultaneous equations
x^= 4 - y
x= y + 2

thank u
AA

Answer by swincher4391(1107)

I'm going to assume that the carat sign was unintentional. Following that assumption:
x = 4-y
x = 2 +y
Add the two equations together
2x = 6 + 0y
x = 3
Plug back in to either equation
3 = 4-y
-1 = -y
y = 1
The solution is (3,1) or x=3 y=1